# Variable aggregation for nonlinear optimization problems

### v2

@@ Abstract
Keywords:
Nonlinear optimization Presolve Graph theory
00footnotetext: We gratefully acknowledge funding from the Los Alamos National Laboratory Laboratory-Directed Research & Development program through the Center for Nonlinear Studies. Approved for unlimited release. LA-UR 25-21592.
@@ 1 Introduction
$\begin{array}[]{c}x^{2}+y^{2}\leq 1\\ y=2z+1\\ \end{array}$
then the expression $12z+$ 1 may be substituted for $y$ in the first constraint to reduce the number of variables and constraints in the optimization problem. We refer to this substitution as variable aggregation and the resulting optimization problem as a reduced-space formulation. In a summary of Gurobi’s gurobimanual pre-solve methods for mixed-integer linear programming, Achterberg et al. achterberg2020 show that variable aggregation can be a source of significant speed-up, but also mention that it can be a source of numerical errors if not implemented carefully. In the context of nonlinear optimization, Bongartz bongartzthesis tests manually-chosen variable aggregations and remarks that “the reduced-space formulations reduced computational times and increased success rates of local solvers” with the caveat that “the smallest possible formulation did not necessarily result in the best computational performance” (see Sections 3.4.3.3 and 3.5 of bongartzthesis ).
@@
Several existing software packages implement automatic variable aggregation for nonlinear optimization models. The AMPL algebraic modeling language ampl scans constraints to identify those written in the form “$y=f(x,\dots)$”. If $y$ is unbounded and has not already been aggregated, the expression $f(x,\dots)$ will be substituted for $y$ everywhere it appears. (This behavior is controlled by the substout option. See fourer1994 and Chapters 14 and 18.2 of ampl for details.) The AIMMS modeling language aimmsmanual and Pyomo modeling environment pyomobook ; pyomopaper also implement presolve options for aggregating variables, but only use equality constraints with at most two variables for these aggregations. Variable aggregation may also be performed within nonlinear optimization solvers. The Knitro knitro and CONOPT conopt solvers implement options for variable aggregation knitromanual ; conoptgamsmanual , but it is not clear what criteria these commercial solvers use to select variables and constraints to eliminate.
@@ 2.1 Nonlinear optimization
We address nonlinear optimization problems in the form given by Equation 1, where $z$ is a variable vector in $\mathbb{R}^{n}$ and $f$, $g$, and $h$ are functions with outputs in $\mathbb{R}$, $\mathbb{R}^{m_{g}}$, and $\mathbb{R}^{m_{h}}$.
$\begin{array}[]{cl}\displaystyle\min_{z}&f(z)\\ \text{s.t.}&g(z)=0\\ &h(z)\leq 0\\ &z^{l}\leq z\leq z^{u}\\ \end{array}$ (1)
Constant vectors $z^{l}$ and $z^{u}$ may have coordinates that are $\infty$- ∞ or +$\infty$+ ∞, respectively, if $z$ is unbounded in some coordinates. Lagrange multipliers for equality, inequality, lower bound, and upper bound constraints are denoted $\lambda$, $\nu$, $\gamma^{l}$, and $\gamma^{u}$. For instance, interior point methods are efficient algorithms for arriving at local solutions of problems in this form biegler2010 . They operate by iteratively solving the linear system given in Equation 2 to compute search directions $(dz,ds,d\lambda,d\nu)$.
$\left[\begin{array}[]{cccc}W+\Sigma&&\nabla g&\nabla h\\ &S^{-1}\Gamma^{s}&&I\\ \nabla g^{T}&&&\\ \nabla h^{T}&I&&\\ \end{array}\right]\left(\begin{array}[]{c}dz\\ ds\\ d\lambda\\ d\nu\\ \end{array}\right)=-\left(\begin{array}[]{c}\nabla f+\nabla g^{T}\lambda+\nabla h^{T}\nu+\left(Z_{u}^{-1}-Z_{l}^{-1}\right)\mu\mathbbm{1}\\ \nu-S^{-1}\mu\mathbbm{1}\\ g\\ h+s\\ \end{array}\right)$ (2)
Here, $W$ is the Hessian of the Lagrangian with respect to $z$ and $\Sigma=Z_{u}^{-1}\Gamma^{u}+Z_{l}^{-1}\Gamma^{l}$. Diagonal matrix $S$ has a diagonal of slack variables $s$, which are introduced along with the bound constraint $s\geq 0$ and multiplier $\gamma^{s}$ to reformulate inequality constraints. $Z_{u}$ and $Z_{l}$ are diagonal matrices of $(z^{u}-z)$ and $(z-z^{l})$, $\Gamma^{s}$, $\Gamma^{u}$, and $\Gamma^{l}$ are diagonal matrices of the corresponding Lagrange multipliers, and $\mu$ a barrier parameter chosen by the algorithm. The matrix on the left-hand-side of this equation is referred to as the Karush-Kuhn-Tucker, or KKT, matrix. Our convention for Jacobian matrices, e.g., $\nabla g$, is that rows correspond to variables while columns correspond to constraints. We note that the transpose of the Jacobian, e.g., $\nabla g^{T}$, has the same sparsity pattern as the incidence matrices we will describe in Section 2.3.
Constructing the linear system in Equation 2 requires the Hessian of the Lagrangian, the Jacobian of constraints, and the gradient of the objective function. If algebraic expression graphs for constraint and objective functions are available, these derivatives are computed by automatic differentiation (AD) of these functions griewank2008 . For most problems, the computational bottleneck of solving Equation 1 with interior point methods is factorization of the KKT matrix. However, for some problems with complicated algebraic expressions, evaluation of the Hessian of the Lagrangian, $W$, is the bottleneck.
@@ 2.2 Algebraic variables and constraints
Section 2.1 considers vector-valued functions of a variable vector $z$. In algebraic modeling language interfaces to nonlinear optimization solvers, however, variables are represented as scalar quantities that participate in algebraic expressions. These algebraic expressions define each coordinate of the functions in Equation 1. An algebraic expression can be represented programmatically as a tree where leaf nodes are (scalar) variables or constants and all other nodes are algebraic operators such as addition ($+$), subtraction (--), or multiplication (*). For example, the algebraic expression tree for the expression $13x^{2}-$ 1 is given in Figure (1).
$\hat{}$$x$22221111
Figure 1: Algebraic expression tree for the expression $13x^{2}-$ 1
@@ 2.4 Variable aggregation
Given an optimization problem in the form of Equation (1), it may be possible to partition the equality constraint functions $g=(\bar{g},\tilde{g})$ and variables $z=(u,v)$ such that the Jacobian $\nabla_{v}\tilde{g}$ is always nonsingular (over a domain of interest). In this case, by the implicit function theorem, there exists a function $v=\tilde{g}_{v}(u)$ that satisfies $\tilde{g}(u,\tilde{g}_{v}(u))=0$. The function $\tilde{g}_{v}$ is exploited in reduced space optimization methods such as parker2022 ; pacaud2022 ; pacaud2024 to eliminate variables and equality constraints (here, $v$ and $\tilde{g}(u,v)=0$) from the optimization problem. The resulting reduced-space optimization problem is given by Equation 3.
$\begin{array}[]{cl}\displaystyle\min_{u}&f(u,\tilde{g}_{v}(u))\\ \text{s.t.}&\bar{g}(u,g_{v}(u))=0\\ &h(u,g_{v}(u))\leq 0\\ &u^{l}\leq u\leq u^{u}\\ &v^{l}\leq g_{v}(u)\leq v^{u}\\ \end{array}$ (3)
@@
Identifying whether a general subset of variables and constraints $v$ and $\tilde{g}(u,v)=0$ is nonsingular over a domain of interest involves solving a global optimization problem, which is beyond the scope of our intended presolve application. However, it is sufficient (but not necessary) for the constraints to have the form given by Lemma 1.
@@
be a subset of constraints in Equation 3, where $\nabla_{v}\tilde{g}^{\mathrm{def},T}$ is strictly lower triangular. Then $\nabla_{v}\tilde{g}^{T}$ is nonsingular everywhere and the subsets of variables and constraints $v$ and $\tilde{g}$ may be eliminated to form a reduced-space optimization problem.
@@
Constraints with this form have the additional advantage that a reduced-space optimization problem may be constructed explicitly by recursively substituting each eliminated variable $v_{i}$ with the corresponding coordinate of the defining function $\tilde{g}^{\rm def}_{i}$. Because of strict lower triangularity of $\tilde{g}^{\rm def}$, the eliminated variable $v_{i}$ does not appear nonlinearly in the equation used to eliminate it, so it is never necessary to solve for $v_{i}$ by an iterative (or implicit) equation solving method.
@@ 3.1 Approximate maximum elimination
The proof is by reduction to the minimum tearing problem. Let $X$ and $C$ be sets of variables and linear constraints with the same cardinalities that admit a perfect matching. It is sufficient to show that identifying a maximum-cardinality matching of $X$ and $C$ with a lower triangular induced submatrix is NP-complete. Identifying this maximum-cardinality matching is equivalent to identifying a the largest square submatrix of $X$ and $C$’s incidence matrix that is lower triangular, which is the tearing problem as described by Carpanzano carpanzano2000order . Carpanzano proves that this problem is NP-complete by reduction to the minimum feedback arc set problem (Problem 8 on Karp’s list of NP-complete problems karp1972reducibility ). Because this problem is a special case of our maximum aggregation set problem, the maximum aggregation set problem is NP-complete as well.
@@ 3.1.1 Greedy heuristic
3.1.2 Matching-based algorithm
@@
      If size of the block is $>1$, the $\tt greedy$ heuristic (Algorithm 1) is used on the block to extract the variable constraint pairs to be added to $\mathcal{M}_{T}$.
@@
The runtime of this algorithm is dominated by either the runtime of computing the maximum linear matching $M_{L}$, or, if $G_{L}$ is small, the runtime of constructing $G_{L}$ in the first place.
@@
2:$G=\text{\tt bipartite\_graph}(X,C)$
3:$G_{L}=\text{\tt linear\_bipartite\_graph}(X,C)$
4:$\mathcal{M}_{L}=\text{\tt maximum\_matching}(G_{L})$
5:$G_{\mathcal{M}}=\text{\tt induced\_subgraph}(G,\mathcal{M}_{L})$
6:$\mathcal{T}=\text{\tt block\_triangularize}(G_{\mathcal{M}},\mathcal{M}_{L})$
7:$\mathcal{M}_{T}=\{\}$
8:for $T\text{~{}\bf in }\mathcal{T}$ do
@@
Let $X$ be a set of variables and $C$ a set of constraints, with bipartite graph $G$ and linear bipartite graph $G_{L}$. Let $n_{\rm match}$ be the cardinality of $\mathcal{M}_{L}$, a maximum matching of the linear bipartite graph. Let $G_{\mathcal{M}}$ be the subgraph of $G$ induced by $\mathcal{M}_{L}$, with a block triangular partition $\mathcal{T}$ of cardinality $n_{\rm block}$. Let $\mathcal{A}$ be the matching returned by Algorithm 2 that is used for aggregation, with cardinality $n_{\rm agg}$. Then
$n_{\rm block}\leq n_{\rm agg}\leq n_{\rm match}$
@@
$n_{\rm agg}\leq\sum_{T\in\mathcal{T}}{\tt size}(T)=n_{\rm match}$
@@
3.2 Fixed-variable aggregation
@@
2:$C^{\prime}=\text{\tt filter\_constraints\_fixed\_variable}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
@@ 3.3 Structure-preserving aggregation
3.3.1 Incidence-preserving aggregation
@@
2:$C^{\prime}=\text{\tt filter\_constraints\_degree2}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
@@
3.3.2 Linearity-preserving elimination
@@
which has linear coefficients of $[1~{}-200]$ (when both variables are brought to the left-hand-side) instead of the original coefficients of $[1~{}-2]$. A similar algorithm is implemented in the Pyomo nl_v2 writer with the linear_presolve=True option.
@@
2:$C^{\prime}=\text{\tt filter\_constraints\_linear\_degree2}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
@@
3.3.3 Equal coefficient degree-2 aggregation
@@
2:$C^{\prime}=\text{\tt filter\_constraints\_equal\_coefficient}(C)$
3:$\mathcal{M}=\text{\tt linear\_matching}(X,C^{\prime})$
@@ 4 Test problems
This section briefly describes the test problems on which we evaluate the methods described in Section 3. To study the effect of aggregation method on solve time, we solve nominal instances of each of the four test problems after applying each aggregation method. To study the affect of aggregation method on convergence reliability, for each of the three test problems that support parameters and each of the aggregation methods, we sample two parameters at eleven uniformly spaced points each (121 total points) and perform a parameter sweep.
@@ 4.1 Distillation column dynamic optimization
Distillation is a widely employed separation technique based on the difference in relative volatility. Dynamic distillation column models play an important role in optimally controlling the distillation column. The dynamic model consists of differential algebraic equations (DAEs) for mass balance that scale with the number of trays, control horizon, and the time discretization along with non-linear constraints linking the mole fractions. The resulting problem is a nonlinear dynamic optimization problem. The full formulation for this model (DIST) may be found in parker2022 .

### v3

@@ 1 Introduction
[…]
then the expression $2z+1$ may be substituted for $y$ in the first constraint to reduce the number of variables and constraints in the optimization problem. We refer to this substitution as variable aggregation and the resulting optimization problem as a reduced-space formulation. In a summary of Gurobi’s gurobimanual pre-solve methods for mixed-integer linear programming, Achterberg et al. achterberg2020 show that variable aggregation can be a source of significant speed-up, but also mention that it can be a source of numerical errors if not implemented carefully. In the context of nonlinear optimization, Bongartz bongartzthesis tests manually-chosen variable aggregations and remarks that “the reduced-space formulations reduced computational times and increased success rates of local solvers” with the caveat that “the smallest possible formulation did not necessarily result in the best computational performance” (see Sections 3.4.3.3 and 3.5 of bongartzthesis).
@@
Several existing software packages implement automatic variable aggregation for nonlinear optimization models. The AMPL algebraic modeling language ampl scans constraints to identify those written in the form $y=f(x,\dots)$”. If $y$ is unbounded and has not already been aggregated, the expression $f(x,\dots)$ will be substituted for $y$ everywhere it appears. (This behavior is controlled by the substout option. See fourer1994 and Chapters 14 and 18.2 of ampl for details.) The AIMMS modeling language aimmsmanual and Pyomo modeling environment pyomobook; pyomopaper also implement presolve options for aggregating variables, but only use equality constraints with at most two variables for these aggregations. Variable aggregation may also be performed within nonlinear optimization solvers. The Knitro knitro and CONOPT conopt solvers implement options for variable aggregation knitromanual; conoptgamsmanual, but it is not clear what criteria these commercial solvers use to select variables and constraints to eliminate.
@@ 2.1 Nonlinear optimization
We address nonlinear optimization problems in the form given by Problem (1), where $z$ is a variable vector in $\mathbb{R}^{n}$ and $f$, $g$, and $h$ are functions with outputs in $\mathbb{R}$, $\mathbb{R}^{m_{g}}$, and $\mathbb{R}^{m_{h}}$.
[…]
Constant vectors $z^{l}$ and $z^{u}$ may have coordinates that are -$\infty$ or $\infty$, respectively, if $z$ is unbounded in some coordinates. Lagrange multipliers for equality, inequality, lower bound, and upper bound constraints are denoted $\lambda$, $\nu$, $\gamma^{l}$, and $\gamma^{u}$. For instance, interior point methods are efficient algorithms for arriving at local solutions of problems in this form biegler2010. They operate by iteratively solving the linear system given in Equation 2 to compute search directions $(dz,ds,d\lambda,d\nu)$.
$\left[\begin{array}[]{cccc}W+\Sigma&&\nabla g&\nabla h\\ &S^{-1}\Gamma^{s}&&I\\ \nabla g^{T}&&&\\ \nabla h^{T}&I&&\\ \end{array}\right]\left(\begin{array}[]{c}dz\\ ds\\ d\lambda\\ d\nu\\ \end{array}\right)=-\left(\begin{array}[]{c}\nabla f+\nabla g^{T}\lambda+\nabla h^{T}\nu+\left(Z_{u}^{-1}-Z_{l}^{-1}\right)\mu\mathbbm{1}\\ \nu-S^{-1}\mu\mathbbm{1}\\ g\\ h+s\\ \end{array}\right)$
Here, $W$ is the Hessian of the Lagrangian with respect to $z$ and $\Sigma=Z_{u}^{-1}\Gamma^{u}+Z_{l}^{-1}\Gamma^{l}$. Diagonal matrix $S$ has a diagonal of slack variables $s$, which are introduced along with the bound constraint $s\geq 0$ and multiplier $\gamma^{s}$ to reformulate inequality constraints. $Z_{u}$ and $Z_{l}$ are diagonal matrices of $(z^{u}-z)$ and $(z-z^{l})$, $\Gamma^{s}$, $\Gamma^{u}$, and $\Gamma^{l}$ are diagonal matrices of the corresponding Lagrange multipliers, and $\mu$ is a barrier parameter chosen by the algorithm. The matrix on the left-hand-side of this equation is referred to as the Karush-Kuhn-Tucker, or KKT, matrix. Our convention for Jacobian matrices, e.g., $\nabla g$, is that rows correspond to variables while columns correspond to constraints. We note that the transpose of the Jacobian, e.g., $\nabla g^{T}$, has the same sparsity pattern as the incidence matrices we will describe in Section 2.3.
Constructing the linear system in Equation 2 requires the Hessian of the Lagrangian, the Jacobian of constraints, and the gradient of the objective function. If algebraic expression graphs for constraint and objective functions are available, these derivatives are computed by automatic differentiation (AD) of these functions griewank2008. For most problems, the computational bottleneck of solving Problem (1) with interior point methods is factorization of the KKT matrix. However, for some problems with complicated algebraic expressions, evaluation of the Hessian of the Lagrangian, $W$, is the bottleneck.
@@ 2.2 Algebraic variables and constraints
Section 2.1 considers vector-valued functions of a variable vector $z$. In algebraic modeling language interfaces to nonlinear optimization solvers, however, variables are represented as scalar quantities that participate in algebraic expressions. These algebraic expressions define each coordinate of the functions in Problem (1). An algebraic expression can be represented programmatically as a tree where leaf nodes are (scalar) variables or constants and all other nodes are algebraic operators such as addition (++), subtraction (-), or multiplication (*). For example, the algebraic expression tree for the expression $3x^{2}-1$ is given in Figure (1).
-$3^\hat{}$xx2211
Figure 1: Algebraic expression tree for the expression $3x^{2}-1$
@@ 2.4 Variable aggregation
Given an optimization problem in the form of Problem (1), it may be possible to partition the equality constraint functions g=(g¯,g~)g=($\bar$$\tilde{g}$) and variables $z=(u,v)$ such that the Jacobian $\nabla_{v}\tilde{g}$ is always nonsingular (over a domain of interest). In this case, by the implicit function theorem, there exists a function $v=\tilde{g}_{v}(u)$ that satisfies $\tilde{g}(u,\tilde{g}_{v}(u))=0$. The function $\tilde{g}_{v}$ is exploited in reduced space optimization methods such as parker2022; pacaud2022; pacaud2024 to eliminate variables and equality constraints (here, $v$ and $\tilde{g}(u,v)=0$) from the optimization problem. The resulting reduced-space optimization problem is given by Problem 3.
[…]
@@
Identifying whether a general subset of variables and constraints $v$ and $\tilde{g}(u,v)=0$ is nonsingular over a domain of interest involves solving a global optimization problem, which is beyond the scope of our intended presolve application. However, it is sufficient (but not necessary) for the constraints to have the form given by Lemma 1, which uses the following definition:
Definition 1
A matrix $A\in\mathbb{R}^{n\times n}$ is strictly lower triangular if $A_{ij}=0$ for all $i\leq j$.
That is, a strictly lower triangular matrix is lower triangular with an empty diagonal.
@@
be a subset of constraints in Problem 3, where $\nabla_{v}\tilde{g}^{\mathrm{def},T}$ is strictly lower triangular. Then $\nabla_{v}\tilde{g}^{T}$ is nonsingular everywhere and the subsets of variables and constraints $v$ and $\tilde{g}$ may be eliminated to form a reduced-space optimization problem.
@@
Constraints with this form have the additional advantage that a reduced-space optimization problem may be constructed explicitly by recursively substituting each eliminated variable $v_{i}$ with the corresponding coordinate of the defining function $\tilde{g}^{\rm def}_{i}$. Because of strict lower triangularity of ${\nabla}_{v}\tilde{g}^{{\rm def}{,T}}$, the eliminated variable $v_{i}$ does not appear nonlinearly in the equation used to eliminate it, so it is never necessary to solve for $v_{i}$ by an iterative (or implicit) equation solving method.
@@ 3.1 Approximate maximum elimination
The proof is by reduction to the minimum tearing problem. Let $X$ and $C$ be sets of variables and linear constraints with the same cardinalities that admit a perfect matching. It is sufficient to show that identifying a maximum-cardinality matching of $X$ and $C$ with a lower triangular induced submatrix is NP-complete. Identifying this maximum-cardinality matching is equivalent to identifying the largest square submatrix of $X$ and $C$’s incidence matrix that is lower triangular, which is the tearing problem as described by Carpanzano carpanzano2000order. Carpanzano proves that this problem is NP-complete by reduction to the minimum feedback arc set problem (Problem 8 on Karp’s list of NP-complete problems karp1972reducibility). Because this problem is a special case of our maximum aggregation set problem, the maximum aggregation set problem is NP-complete as well.
@@ 3.1.1 Greedy heuristic (GR)
3.1.2 Matching-based algorithm (LM)
@@
      If size of the block is >1>1, the $\tt greedy$ heuristic (Algorithm 1) is used on the block to extract the variable constraint pairs to be added to $\mathcal{M}_{T}$.
@@
The runtime of this algorithm is dominated by either the runtime of computing the maximum linear matching ${\mathcal{M}}_{L}$, or, if $G_{L}$ is small, the runtime of constructing $G_{L}$ in the first place.
@@
[…]
@@
Let $X$ be a set of variables and $C$ a set of constraints, with bipartite graph $G$ and linear bipartite graph $G_{L}$. Let $n_{\rm match}$ be the cardinality of $\mathcal{M}_{L}$, a maximum matching of the linear bipartite graph. Let $G_{\mathcal{M}}$ be the subgraph of $G$ induced by $\mathcal{M}_{L}$, with a block triangular partition $\mathcal{T}$ of cardinality $n_{\rm block}$. Let $\mathcal{A}$ be the matching returned by Algorithm 2 that is used for aggregation, with cardinality $n_{\rm agg}$. Then,
$n_{\rm block}\leq n_{\rm agg}\leq n_{\rm match}{.}$
@@
$n_{\rm agg}\leq\sum_{T\in\mathcal{T}}{\tt size}(T)=n_{\rm match}{.}$
@@
3.2 Fixed-variable aggregation (LD1)
@@
[…]
@@ 3.3 Structure-preserving aggregation
3.3.1 Incidence-preserving aggregation (D2)
@@
[…]
@@
3.3.2 Linearity-preserving elimination (LD2)
@@
which has linear coefficients of [1200][1~-200] (when both variables are brought to the left-hand-side) instead of the original coefficients of [12][1~-2]. A similar algorithm is implemented in the Pyomo nl_v2 writer with the linear_presolve=True option.
@@
[…]
@@
3.3.3 Equal coefficient degree-2 aggregation (ECD2)
@@
[…]
@@ 4 Test problems
This section briefly describes the test problems on which we evaluate the methods described in Section 3. To study the effect of aggregation method on solve time, we solve nominal instances of each of the four test problems after applying each aggregation method. To study the effect of aggregation method on convergence reliability, for each of the three test problems that support parameters and each of the aggregation methods, we sample two parameters at eleven uniformly spaced points each (121 total points) and perform a parameter sweep.
@@ 4.1 Distillation column dynamic optimization
Distillation is a widely employed separation technique based on the difference in relative volatility. Dynamic distillation column models play an important role in optimally controlling the distillation column. The dynamic model consists of differential algebraic equations (DAEs) for mass balance that scale with the number of trays, control horizon, and the time discretization along with non-linear constraints linking the mole fractions. The resulting problem is a nonlinear dynamic optimization problem containing bilinear terms and rational nonlinearities. The full formulation for this model (DIST) may be found in parker2022.